a grocer needs to mix raisins at $2.00 per lbs with granola @ $3.25 per lbs to obtain 10 lbs of mixture that cost $2.50 per lb. how many pounds of each must be used?
Answer by nerdybill(7384)

a grocer needs to mix raisins at $2.00 per lbs with granola @ $3.25 per lbs to obtain 10 lbs of mixture that cost $2.50 per lb. how many pounds of each must be used?
.
Let x = lbs of raisins
10-x = lbs of granola
.
2x + 3.25(10-x) = 2.50(10)
2x + 32.5 - 3.25x = 25
32.5 - 1.25x = 25
32.5 = 25 + 1.25x
7.5 = 1.25x
6 lbs = x (raisins)
.
10-x = 10-6 = 4 lbs (granola)